OpenStudy (awy):

Simplify completely the trigonometric expression. (1-sinx/cosx) + (cosx/1-sinx)

OpenStudy (anonymous):

\[\frac{1-\sin x}{\cos x} + \frac{\cos x}{1-\sin x}\]?

OpenStudy (awy):

yes, trying to figure out how to simplify that.

OpenStudy (anonymous):

(1-sinx)^2 + cos^2x -------------------------- cosx(1-sinx) 1+sin^2x+cos^2x-2sinx -------------------------------- cosx(1-sinx) 2-2sinx ---------- cosx(1-sinx) 2(1-sinx) ------------ cosx(1-sinx) 2 --- cosx

OpenStudy (awy):

aww yea, k, i got that answer too. Thx.